### v2

# Small Models Struggle to Learn from Strong Reasoners

Abstract
Large language models (LLMs) excel in complex reasoning tasks, and distilling their reasoning capabilities into smaller models has shown promise. However, we uncover an interesting phenomenon, which we term the Small Model Learnability Gap: small models (3B parameters) do not consistently benefit from long chain-of-thought (CoT) reasoning or distillation from larger models. Instead, they perform better when fine-tuned on shorter, simpler reasoning chains that better align with their intrinsic learning capacity. To address this, we propose Mix Distillation, a simple yet effective strategy that balances reasoning complexity by combining long and short CoT examples or reasoning from both larger and smaller models. Our experiments demonstrate that Mix Distillation significantly improves small model reasoning performance compared to training on either data alone. These findings highlight the limitations of direct strong model distillation and underscore the importance of adapting reasoning complexity for effective reasoning capability transfer.
Small Models Struggle to Learn from Strong Reasoners
Yuetai Li Xiang Yue Zhangchen Xu Fengqing Jiang Luyao Niu Bill Yuchen Lin Bhaskar Ramasubramanian Radha Poovendran University of Washington Carnegie Mellon University Western Washington University {yuetaili,zxu9,fqjiang,luyaoniu,byuchenl,rp3}@uw.edu, xyue2@andrew.cmu.edu, ramasub@wwu.edu Huggingface: https://huggingface.co/UWNSL Project Page: https://small-model-gap.github.io/
1 Introduction
Large language models (LLMs) (Anthropic, 2023; Brown et al., 2020; OpenAI, 2023; Touvron et al., 2023a) have demonstrated remarkable performance in complex reasoning tasks, enabling advancements in mathematical problem-solving, logical inference, and structured decision-making (Cobbe et al., 2021; Shao et al., 2024; Yang et al., 2024). A key advancement in improving LLM complex reasoning capability is the chain-of-thought (CoT) prompting. This technique decomposes complex problems into intermediate reasoning steps, enhancing both performance and interpretability. (Wei et al., 2023).
However, the high computational cost of LLMs hinders their deployment on resource-constrained devices, motivating the development of smaller models that offer similar capabilities at reduced cost. A widely adopted strategy to achieve this is distillation (Agarwal et al., 2024; Hinton et al., 2015; Kim et al., 2024a), where CoT sequences generated by a strong teacher model are used to fine-tune a weaker student model. Naturally, one might expect that distilling CoT sequences from stronger models would consistently improve small models’ complex reasoning capabilities (Agarwal et al., 2024; DeepSeek-AI et al., 2024; Min et al., 2024; Tunstall et al., 2023).
However, we reveal an interesting phenomenon, which we term the Small Model Learnability Gap (Fig. 1): small models do not consistently benefit from the complex reasoning sequences provided by strong teachers, such as long CoT reasoning or distillation from large models. In our experiments, we observe that when small models are exposed to long and intricate reasoning traces, they struggle to internalize the multi-step logic due to their constrained ability. Instead, small models perform better when fine-tuned on shorter, simpler reasoning chains that align more closely with their intrinsic learning capacity. This suggests that small models struggle to process overly elaborate reasoning traces or adapt to the distribution shifts introduced by stronger teachers, ultimately limiting their ability to generalize effectively.
To address the challenge described above, we propose Mix Distillation, a simple yet effective approach that balances reasoning complexity by blending different types of reasoning traces. Specifically, our method comprises two configurations: (1) Mix-Long – A combination of long and short CoT examples, ensuring that small models are exposed to both detailed and concise reasoning steps. (2) Mix-Large – A mixture of responses from both larger and smaller models, allowing small models to learn from reasoning chains that are better suited to their capacity.
Our experiments demonstrate that Mix Distillation consistently improves small model reasoning performance compared to standard distillation. For instance, Qwen2.5-3B-Instruct improves by more than 8 points on MATH and AMC using Mix-Long, compared to direct training on long CoT data. Qwen2.5-3B-Instruct gains more than 7 points on MATH, AIME and AMC using Mix-Large compared with training on large teacher CoT data.
These findings highlight a fundamental limitation of direct strong model distillation and emphasize the importance of adapting reasoning complexity for effective knowledge transfer. By carefully designing distillation strategies, we provide new insights into overcoming the constraints of small model learning, making them more effective at reasoning-intensive tasks.
2 Preliminaries
2.1 Notation
Let represent an input sequence (e.g., a prompt), and be the corresponding output sequence. We consider a LLM parameterized by , which predicts the next token following a conditional distribution . We denote by the subset of tokens in the generated output that encodes a chain-of-thought, often serving as a reasoning trace or explanatory sequence.
Throughout this work, we use the term short CoT, to describe concise reasoning paths to arrive at solutions (Min et al., 2024; Yeo et al., 2025) and long CoT to describe an extended reasoning sequence that is not only longer but also demonstrates more complex reflective thoughts (Qwen, 2024b; Yeo et al., 2025). Additionally, we use the term large teacher CoT to refer to the reasoning trace generated by a larger teacher model, and the term small teacher CoT for the reasoning steps produced by a smaller teacher model. Please see Appendix D for more examples.
2.2 Supervised Fine-Tuning (SFT)
Supervised fine-tuning (SFT) is widely adopted to enhance reasoning capabilities of LLMs on a dataset , where can be short CoT, long CoT, strong model CoT or weak model CoT sequences. The SFT process updates the parameters of a language model by minimization the negative log-likelihood loss over the instruction dataset .
3 Small Model Learnability Gap
In this section, we fine-tune student models using different CoT data. We then reveal the small model learnability gap given the performance of fine-tuned models.
3.1 Experiment Setup
Datasets.
We use the 7,500 prompt set of MATH (Hendrycks et al., 2021). This dataset encompasses seven math topics such as advanced calculus, geometry, and linear algebra.
Student models.
Our study considers ten student models from the Qwen (Qwen, 2024a) and Llama (Meta, 2024a, b) model families of varying sizes. These models include the Instruct version of Qwen2.5-0.5B, Qwen2.5-1.5B, Qwen2.5-3B, Qwen2.5-7B, Qwen2.5-14B, and Qwen2.5-32B, and the Instruct version of Llama3.2-1B, Llama3.2-3B, Llama3.1-8B, and Llama3.3-70B. A comprehensive overview of the student models is presented in Table 4 of Appendix A.
Teacher models.
To compare long CoT with short CoT, we use QwQ-32B-Preview (Qwen, 2024b) to generate long CoT sequences and Qwen2.5-32B-Instruct as the response generator for short CoT. Within each model family, we designate the larger scale model as the large teacher and the smaller scale model as the small teacher. This includes Qwen2.5-72B-Instruct vs Qwen2.5-3B-Instruct, Llama3.1-70B-Instruct vs Llama3.1-8B-Instruct, and Gemma2-27B-it vs Gemma2-9B-it.
Evaluation Benchmarks.
We evaluate the reasoning capability of fine-tuned student models on a set of commonly used benchmarks, including MATH (Hendrycks et al., 2021), GSM8K (Cobbe et al., 2021), AMC 2023, AIME 2024, and the English math subset of OlympiadBench (He et al., 2024). These benchmarks span a wide range of challenge levels, from elementary mathematics to advanced competition problems. We define the student model performance as the average score on five benchmarks. Unless otherwise specified, all fine-tuned models are evaluated in a zero-shot setting using greedy decoding. We set the maximum generation tokens as 16k. Please see Appendix A for detailed experimental setup.
We define the following performance scores:
: Performance score of a student model fine-tuned on long CoT data.
: Performance score of a student model fine-tuned on short CoT data.
: Performance score of a student model fine-tuned on CoT from a larger teacher.
: Performance score of a student model fine-tuned on CoT from a smaller teacher.
Training Setup.
Teacher models generate responses by rejection sampling (Dong et al., 2023; Gulcehre et al., 2023; Tong et al., 2024; Yuan et al., 2023; Yue et al., 2023; Zelikman et al., 2022) By default, teacher models employ greedy decoding. By combining the math problem instructions with corresponding solutions generated by teacher models, we construct problem-solution pairs to fine-tune student models. We train the models using the LLaMA-Factory framework (Zheng et al., 2024). For student models of scale less than 14B, we use full-parameter SFT and implement a cosine learning rate schedule with a maximum learning rate of to fine-tune student models for two epochs (Touvron et al., 2023b). For student models larger than 14B, we adopt LoRA fine-tuning with a learning rate of for two epochs. Detailed hyperparameters and information about the experimental platform are provided in Appendix A.
3.2 Long CoT Gap
This section evaluates the reasoning capabilities of student models fine-tuned over long CoT data and short CoT data. We quantify the performance difference between long and short CoT data using long CoT gap , defined as:
Figure 2 provides a comprehensive overview of the long CoT gap across different student models. The detailed benchmark scores on MATH, GSM8K, AIME, AMC, and OlympiadBench are deferred to Table 7 in Appendix B. We report the following key takeaways.
We observe that long CoT is more effective for larger models, consistently leading to improved performance across most math benchmarks. For example, the student model Qwen2.5-32B-Instruct improves about 15 points across all math metrics on average.
However, long CoT data is not effective for smaller models, yielding significantly less improvement compared to short CoT. On the MATH and AMC benchmarks, student model Qwen2.5-1.5B-Instruct performs over 10 points lower when fine-tuned with long CoT data. This shows that smaller models may not be able to effectively learn and utilize the long CoT paradigm. Please see more attribution analysis in Section 3.4.
Student Model | Better? | |||
Qwen2.5-0.5B | 14.8 | 19.5 | -4.7 | Short |
Qwen2.5-1.5B | 27.0 | 34.2 | -7.1 | Short |
Qwen2.5-3B | 40.3 | 43.4 | -3.1 | Short |
Qwen2.5-7B | 48.9 | 47.2 | 1.7 | Long |
Qwen2.5-14B | 59.2 | 54.3 | 4.9 | Long |
Qwen2.5-32B | 73.0 | 59.3 | 13.7 | Long |
Llama-3.2-1B | 15.8 | 19.5 | -3.7 | Short |
Llama-3.2-3B | 32.5 | 33.1 | -0.6 | Short |
Llama-3.1-8B | 35.2 | 31.5 | 3.7 | Long |
Llama-3.3-70B | 58.2 | 54.3 | 3.8 | Long |
Student Model | Better? | |||
Qwen2.5-0.5B | 16.9 | 20.4 | -3.5 | Weak |
Qwen2.5-1.5B | 32.2 | 33.0 | -0.8 | Weak |
Qwen2.5-3B | 39.7 | 39.4 | 0.3 | Strong |
Qwen2.5-7B | 48.9 | 42.3 | 6.6 | Strong |
Qwen2.5-14B | 52.9 | 49.9 | 3.0 | Strong |
Qwen2.5-32B | 59.5 | 53.0 | 6.5 | Strong |
Llama-3.2-1B | 16.5 | 18.5 | -1.9 | Weak |
Llama-3.2-3B | 32.8 | 31.2 | 1.6 | Strong |
Llama-3.1-8B | 25.6 | 25.1 | 0.5 | Strong |
Llama-3.3-70B | 57.6 | 53.3 | 4.3 | Strong |
3.3 Large Teacher CoT Gap
We investigate how effective small models may learn from large teacher and small teachers. We define a large teacher CoT gap as:
Figure 3 provides a comprehensive comparison of the incurred by all student models. The detailed benchmark scores of MATH, GSM8K, AIME, AMC and OlympiadBench are deferred to Table 8 in Appendix B. More experimental results of different teacher models, including Llama3.1-70B vs Llama3.1-8B and Gemma2-27B vs Gemma2-9B are in Table 9 of Appendix B.
We observe that larger student models learn effectively from large teacher CoT. For example, Qwen2.5-7B-Instruct and Qwen2.5-32B-Instruct student models improve over 5 points on average, with Qwen2.5-32B-Instruct achieving more than a 15 point increase on the AIMC benchmark. However, smaller models do not learn effectively from large teacher models such as Qwen2.5-72B-Instruct. Instead, small teacher models such as Qwen2.5-3B-Instruct may serve as better teacher models for small student models. For instance, the performance of Qwen2.5-0.5B-Instruct degrades by more than 10 points on the AMC benchmark.
Note that prior studies (Kim et al., 2024b) also demonstrated that stronger models are not necessarily stronger teachers, emphasizing response generator and teacher-side factors. Our work differs in that we attribute this phenomenon primarily to the size of the student model.
Student Model | Distillation Method | MATH | AMC | GSM8k | Olympiad Bench | AIME | Average |
Qwen2.5-3B | Long CoT | 56.2 | 37.5 | 80.0 | 24.4 | 3.3 | 40.3 |
Short CoT | 61.0 | 37.5 | 82.0 | 26.4 | 10.0 | 43.4 | |
Large Model CoT | 57.5 | 35.0 | 80.0 | 25.9 | 0.0 | 39.7 | |
Small Model CoT | 60.3 | 27.5 | 79.5 | 26.4 | 3.3 | 39.4 | |
Deepseek-R1-32B (Long CoT) | 50.7 | 20.0 | 81.2 | 15.7 | 0.0 | 33.5 | |
Ours | |||||||
Mix-Long | 64.7 | 45.0 | 81.4 | 28.6 | 10.0 | 45.9 | |
Mix-Large | 65.8 | 42.5 | 81.7 | 29.0 | 10.0 | 45.8 | |
Llama3.2-3B | Long CoT | 48.7 | 17.5 | 75.1 | 17.6 | 3.3 | 32.5 |
Short CoT | 50.9 | 15.0 | 77.5 | 18.7 | 3.3 | 33.1 | |
Large Model CoT | 47.4 | 25.0 | 71.2 | 16.9 | 3.3 | 32.8 | |
Small Model CoT | 47.9 | 17.5 | 74.1 | 16.4 | 3.3 | 31.2 | |
Deepseek-R1-32B (Long CoT) | 48.5 | 17.5 | 77.7 | 16.1 | 6.7 | 33.3 | |
Ours | |||||||
Mix-Long | 53.0 | 22.5 | 79.4 | 17.2 | 3.3 | 35.1 | |
Mix-Large | 51.8 | 25.0 | 76.3 | 17.2 | 3.3 | 34.7 |
3.4 Analysis of Small Model Learnability Gap
Domain knowledge affects learnability gap.
We observe that math expert models, in spite of small model size, exhibit a smaller learnability gap for both long CoT and large teacher CoT data compared to general models in Figure 4. Specifically, we compare the learnability gaps between the student models Qwen2.5-Math-1.5B-Instruct and Qwen2.5-1.5B-Instruct. Our findings show that the long CoT gap of the small math expert model is significantly smaller than that of general small models. Furthermore, the performance improvement of Qwen2.5-Math-1.5B when fined-tuned with large teacher CoT exceeds that of Qwen2.5-1.5B, suggesting that math expert models benefit more substantially from large teacher CoT. We conjecture that a key factor leading to the small model learnability gap is the limited in-domain knowledge of small student models. We summarize this observation in the following takeaway.
Base models exhibit a more significant learnability gap.
We observe that base models generally exhibit a more significant learnability gap than Instruct models in Figure 5. This suggests that it is more challenging for small base models to effectively learn from long CoT data or large teacher CoT.
Speaking styles shift.
We adopt the method from (Lin et al., 2023) to evaluate the rank shift of each token before and after fine-tuning on long CoT and Large teacher CoT data. This allows us to compare the token distribution shifts induced by the fine-tuning process. We then annotate the tokens that exhibit the largest rank shifts as the most shifted tokens. Our analysis reveals that these tokens are predominantly associated with expressive and stylistic elements, such as “wait”, “But”, and “Let”. Please see Appendix C for more details.
4 Mix Distillation: Bridge Small Model Learnability Gap
This section presents our Mix Distillation approach to bridge the small model learnability gap.
4.1 Mix Distillation
We propose Mix Distillation to address the learnability gap observed in small models. This approach blends easier-to-learn data with more challenging data for small models, thereby leveraging the strengths of both.
Our insight is that small models tend to perform better on data that closely matches their inherent distribution (such as short CoT or small teacher CoT), while they struggle with data that exhibits greater distribution shifts. The token distribution of the mixed long CoT and large teacher CoT data may become closer to that of small models’ inherent distribution, thereby enabling them to learn more effectively from challenging datasets.
We propose Mix-Long, which combines long CoT and short CoT data with a weight of long CoT and short CoT . Similarly, we proposed Mix-Large, which combines large teacher CoT with a weight of and small teacher CoT with a weight of .
4.2 Experiment Results
We use Qwen2.5-3B-Instruct as the student model and MATH (7.5k) as the training set. We distill different teacher models to generate responses as the baseline. They include QwQ-32B (long CoT), Qwen2.5-32B (short CoT), Qwen2.5-72B (large teacher CoT), Qwen2.5-3B (small teacher CoT). We add Deepseek-R1-32B (DeepSeek-AI, 2025) as the teacher model to generate another set of long CoT data as baseline. We set in both configurations of Mix-Long and Mix-Large.
Experimental results demonstrate that both Mix-Long and Mix-Large surpass baselines in most evaluation metrics. We show that the small student model could achieve improved performance by Mix Distillation compared to training on a single dataset. For instance, Qwen2.5-3B-Instruct improves by more than 8 points on MATH and AMC using Mix-Long, compared to direct training on long CoT data. It also shows a more than 7-point gain on MATH, AIME and AMC for Qwen2.5-3B-Instruct by Mix-Large compared with training on large teacher CoT data. This implies that it is easier for small student models to learn from datasets generated by Mix Distillation.
Figure 6 shows the average performance when taking different mix weight of long CoT data or large teacher CoT. We choose Qwen2.5-3B-Instruct as the student model and find that a weight of 0.2 achieves the highest average performance across five benchmarks for both Mix-Long and Mix-Large.
Interestingly, we find that after mixing long CoT and short CoT data, the small student model’s output incorporates characteristics of long CoT, such as a branching process, while maintaining a reduced token length and avoiding overly elaborate thinking. This is illustrated in Figure 7. We observed that the small student model fine-tuned on long CoT data becomes overwhelmed by repeated thoughts and fails to stop, whereas the model fine-tuned on short CoT data produces incorrect answers. In contrast, our proposed Mix-Long, which incorporates branching elements (e.g., the use of “Alternatively”), delivers the correct answer. Additionally, the average token lengths of responses generated by long CoT, short CoT, and Mix-Long are 3384.7, 575.7, and 1248.9, respectively. We suggest that mixing long CoT and short CoT data is a practical approach to achieving a balanced CoT length, thereby enhancing the reasoning capabilities of small student models.
5 Related Work
5.1 Chain-of-Thought
Early research on CoT primarily focused on short CoT, where models produce succinct reasoning paths to reach a solution (Lambert et al., 2025; Longpre et al., 2023; Wei et al., 2023; Yu et al., 2024). Recently, researchers have turned to long CoT prompting, which encourages the generation of extended and detailed reasoning chains (DeepSeek-AI, 2025; Hou et al., 2025; Kimi Team, 2025; NovaSky, 2025; OpenAI, 2024; Pan et al., 2025; Zeng et al., 2025). The model systematically explores multiple paths (branching) and reverts to earlier points if a particular path proves wrong (backtracking). Although several studies have investigated methods such as distillation and reinforcement learning to integrate long CoT capabilities into LLMs, these efforts have predominantly concentrated on large models. In contrast, our work specifically targets the challenges associated with training smaller models.
5.2 Synthetic Reasoning Data
Although human-crafted reasoning datasets have been used to enhance LLM reasoning capabilities Hendrycks et al. (2021); LI et al. (2024), their development is both time-consuming and labor-intensive. Recent advancements have streamlined this process by generating instructions or responses directly from LLMs (Hui et al., 2024; Toshniwal et al., 2024; Xu et al., 2024; Yue et al., 2023; Zhang et al., 2025) or extracting data directly from web (Paster et al., 2023; Yue et al., 2024), yielding more detailed and diverse chain-of-thought reasoning pathways. Recent study has investigated the impact of various response generators (Kim et al., 2024b), suggesting that in the domains of instruction following and reasoning, responses from stronger teacher models do not necessarily produce the most effective learning effects for student models. However, these investigations have not recognized student model size as a critical factor influencing this phenomenon, nor have they performed the more attribution and mitigation analyses as in this paper.
6 Conclusion and Future Work
In this paper, we show that long CoT data and large model responses were not uniformly beneficial for small student models. We found that small models may perform better when fine-tuned with short CoT and small model CoT. We termed this challenge as the Small Model Learnability Gap. The reason behind it may be that small student models excel on data that closely match their inherent distribution but struggle with significant distribution shifts. To bridge the gap, we introduced Mix Distillation, including Mix-Long, which combined long CoT and short CoT data in a ratio, and Mix-Large, which integrated large and small teacher CoT. Experimental results showed that both Mix-Long and Mix-Large outperform baselines across most evaluation metrics, which implied mix distillation outperforms training on a single data distribution. This paper provided practical insights for optimizing post-training strategies to enhance small language model reasoning capability.
We will explore several promising directions as future work. First, we will refine mix distillation by optimally combining diverse data sources and proposing more fine-grained mixing algorithms to boost reasoning capabilities. Second, we propose to study how strong reasoning teachers can generate data that is better suited for tuning small student models, thereby facilitating more effective knowledge transfer. Third, we will conduct further theoretical and model interpolability studies on the small model learnability gap. Lastly, we will investigate which SFT methods yield the best initial policies for subsequent RL procedure, ultimately enhancing overall model performance.
Acknowledgment
This work is partially supported by the Air Force Office of Scientific Research (AFOSR) under grant FA9550-23-1-0208, the Office of Naval Research (ONR) under grant N0014-23-1-2386, and the National Science Foundation (NSF) AI Institute for Agent-based Cyber Threat Intelligence and Operation (ACTION) under grant IIS 2229876.
This work is supported in part by funds provided by the National Science Foundation, Department of Homeland Security, and IBM. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the NSF or its federal agency and industry partners.
Limitations
While our study provides valuable insights into the understanding of small model learnability gap in math reasoning, we acknowledge that our research primarily focuses on this specific domain and does not evaluate other crucial skills such as instruction following, code generation, or multi-modal understanding. We also did not investigate the impact of fine-grained variations in pre-training data composition on the small model learnability gap. A more detailed analysis of how different pre-training data sources and their proportions affect learning outcomes could offer valuable insights into optimizing data selection strategies for mitigating this gap.
Ethical Statement
This paper focuses on the evaluation and enhancement of reasoning capabilities in small language models through distillation techniques. The dataset and benchmarks used in our experiments are publicly available. We do not introduce or endorse any applications that could cause harm or be misused. This paper does not present any ethical concerns.
Appendix A Detailed Experimental Setups
Category | Models |
Teacher Models | |
Long CoT vs | QwQ-32B-Preview vs |
ShortCoT | Qwen2.5-32B-Instruct |
Large Teacher vs | |
Small Teacher | |
Qwen Family | Qwen2.5-72B-Instruct vs |
Qwen2.5-3B-Instruct | |
Llama Family | Llama3.1-70B-Instruct vs |
Llama3.1-8B-Instruct | |
Gemma Family | Gemma2-27B-it vs |
Gemma2-9B-it | |
Student Models | |
Qwen Family | Qwen2.5-0.5B-Instruct, |
Qwen2.5-1.5B-Instruct, | |
Qwen2.5-3B-Instruct, | |
Qwen2.5-7B-Instruct, | |
Qwen2.5-14B-Instruct, | |
Qwen2.5-32B-Instruct | |
Llama Family | Llama3.2-1B-Instruct, |
Llama3.2-3B-Instruct, | |
Llama3.1-8B-Instruct, | |
Llama3.3-70B-Instruct |
A.1 Models
Table 4 presents a comprehensive overview of student and teacher models used in our paper.
A.2 Training Setup
Our model training is conducted using LLaMA-Factory (Zheng et al., 2024), on a server with four NVIDIA A100-SXM4-80GB GPUs, an AMD EPYC 7763 64-Core Processor, and 512 GB of RAM. We use full parameter fine-tuning on student models less than 14B parameters. When the student model is larger than 14B, we use LoRA fine-tuning Hu et al. (2021). Table 5 and Table 6 list hyper-parameters for full parameter fine-tuning and LoRA fine-tuning respectively.
Hyper-parameter | Value |
Learning Rate | |
Number of Epochs | |
Number of Devices | |
Per-device Batch Size | |
Optimizer | Adamw |
Learning Rate Scheduler | cosine |
Max Sequence Length |
Hyper-parameter | Value |
Learning Rate | |
Number of Epochs | |
Number of Devices | |
Per-device Batch Size | |
Lora Target | full |
Learning Rate Scheduler | cosine |
Warmup Ratio | |
Max Sequence Length |
Teacher models generate responses by rejection sampling (Zelikman et al., 2022; Tong et al., 2024; Yue et al., 2023; Singh et al., 2024; Gulcehre et al., 2023; Yuan et al., 2023; Dong et al., 2023). The prompt used to make teacher models generate responses to the math questions is shown as follows. By default, teacher models employ greedy decoding. By combining the math problem instructions with corresponding solutions generated by teacher models, we construct problem-solution pairs to fine-tune student models. We perform pairwise comparisons of solutions generated by different teacher models and filter out problem-solution pairs that are correct for both models to fine-tune student models.
A.3 Evaluation Setup
We evaluate the reasoning capability of fine-tuned student models on a set of commonly used benchmarks, including MATH (Hendrycks et al., 2021), GSM8K (Cobbe et al., 2021), AMC 2023, AIME 2024, and the English math subset of OlympiadBench (He et al., 2024).
Unless otherwise specified, all fine-tuned models are evaluated in a zero-shot setting using greedy decoding. We set the maximum generation tokens as 16k. The evaluation prompt is shown below.
After extracting the final answer of the evaluated model, we first employ exact matching to determine the correctness of the answer. If the answer is incorrect, we use Qwen-32B-Instruct as a judge to compare the extracted final answers against that of the ground truth. The prompt is shown below.
Appendix B More Experiments Results
In this section we present additional experiment results of long CoT gap and large teacher CoT gap.
B.1 Long CoT Gap: Additional Results
Table 7 shows the detailed performance scores and gap of each benchmark for different student models fine-tuned on long CoT and short CoT. QwQ-32B-Preview is chosen to generate long CoT and awhile Qwen-2.5-32B-Instruct is chosen to generate short CoT. We observe that small student models tend to benefit more from short CoT, while large student models gain greater advantages from long CoT.
MATH | GSM8K | AIME | AMC | Olympiad | Average | |||||||||||
Model | ||||||||||||||||
Llama-3.2-1B | 28.6 | 33.4 | -4.78 | 42.3 | 49.2 | -6.90 | 0.00 | 0.00 | 0.00 | 2.50 | 7.50 | -5.00 | 5.48 | 7.40 | -1.92 | -3.72 |
Llama-3.2-3B | 48.7 | 50.9 | -2.14 | 75.1 | 77.5 | -2.42 | 3.33 | 3.33 | 0.00 | 17.5 | 15.0 | 2.50 | 17.6 | 18.7 | -1.04 | -0.619 |
Llama-3.1-8B | 50.0 | 44.6 | 5.36 | 81.4 | 75.5 | 5.84 | 0.00 | 0.00 | 0.00 | 27.5 | 22.5 | 5.00 | 17.3 | 14.8 | 2.52 | 3.74 |
Llama-3.3-70B | 75.3 | 74.9 | 0.340 | 92.7 | 91.2 | 1.44 | 26.7 | 13.3 | 13.3 | 55.0 | 52.5 | 2.50 | 41.3 | 39.7 | 1.63 | 3.85 |
Qwen2.5-0.5B | 23.0 | 31.5 | -8.44 | 39.5 | 45.3 | -5.84 | 0.00 | 0.00 | 0.00 | 7.50 | 15.0 | -7.50 | 4.00 | 5.93 | -1.93 | -4.74 |
Qwen2.5-1.5B | 41.6 | 52.3 | -10.7 | 63.8 | 71.7 | -7.89 | 0.00 | 0.00 | 0.00 | 17.5 | 27.5 | -10.0 | 12.3 | 19.4 | -7.11 | -7.13 |
Qwen2.5-3B | 56.2 | 61.0 | -4.84 | 80.0 | 82.0 | -1.98 | 3.33 | 10.0 | -6.67 | 37.5 | 37.5 | 0.00 | 24.4 | 26.4 | -1.93 | -3.08 |
Qwen2.5-7B | 68.2 | 67.8 | 0.460 | 86.2 | 85.7 | 0.560 | 13.3 | 6.67 | 6.67 | 40.0 | 40.0 | 0.00 | 36.6 | 35.7 | 0.889 | 1.72 |
Qwen2.5-14B | 78.3 | 76.2 | 2.04 | 93.3 | 92.5 | 0.760 | 20.0 | 6.67 | 13.3 | 60.0 | 55.0 | 5.00 | 44.4 | 40.9 | 3.56 | 4.94 |
Qwen2.5-32B | 84.8 | 82.3 | 2.44 | 94.9 | 94.3 | 0.610 | 40.0 | 10.0 | 30.0 | 85.0 | 62.5 | 22.5 | 60.4 | 47.3 | 13.2 | 13.7 |
B.2 Large Teacher CoT Gap: Additional Results
Table 8 shows the detailed performance scores and gap of each benchmark for different student models distilled from large teacher and small teacher. We summarize the performance of 10 student models from the Llama and Qwen families across various model sizes. Qwen-2.5-72B-Instruct is chosen as the large teacher while Qwen-2.5-3B-Instruct is chosen as the small teacher. The results are shown in Table 8. Our findings indicate that small student models may experience degraded performance when distilled from a large teacher compared to a small teacher, whereas larger student models benefit more from distilling a large teacher.
Table 9 shows more experiment results for teacher models in different model families, including Gemma-27B-it vs Gemma-9B-it and Llama3.1-72B-Instruct vs Llama3.1-8B-Instruct.
MATH | GSM8k | AIME | AMC | Olympiad | Average | |||||||||||
Model | ||||||||||||||||
Llama-3.2-1B | 29.8 | 29.6 | 0.160 | 44.4 | 47.5 | -3.18 | 0.00 | 0.00 | 0.00 | 2.50 | 7.50 | -5.00 | 6.07 | 7.70 | -1.63 | -1.93 |
Llama-3.2-3B | 47.4 | 47.9 | -0.500 | 71.2 | 74.1 | -2.88 | 3.33 | 0.00 | 3.33 | 25.0 | 17.5 | 7.50 | 16.9 | 16.4 | 0.445 | 1.58 |
Llama-3.2-8B | 37.6 | 37.6 | -0.040 | 67.0 | 69.2 | -2.20 | 6.67 | 0.00 | 6.67 | 7.50 | 7.50 | 0.00 | 9.19 | 11.0 | -1.78 | 0.530 |
Llama-3.2-70B | 74.5 | 72.2 | 2.28 | 92.0 | 92.2 | -0.152 | 16.7 | 16.7 | 0.00 | 67.5 | 50.0 | 17.5 | 37.3 | 35.7 | 1.63 | 4.25 |
Qwen2.5-0.5B | 30.0 | 31.0 | -0.920 | 43.1 | 45.4 | -2.35 | 0.00 | 0.00 | 0.00 | 5.00 | 17.5 | -12.5 | 6.52 | 8.30 | -1.78 | -3.51 |
Qwen2.5-1.5B | 50.3 | 50.7 | -0.440 | 70.6 | 71.0 | -0.455 | 0.00 | 3.33 | -3.33 | 22.5 | 20.0 | 2.50 | 17.8 | 20.0 | -2.22 | -0.790 |
Qwen2.5-3B | 57.5 | 60.3 | -2.82 | 79.9 | 79.5 | 0.379 | 0.00 | 3.33 | -3.33 | 35.0 | 27.5 | 7.50 | 25.9 | 26.4 | -0.444 | 0.256 |
Qwen2.5-7B | 71.3 | 63.6 | 7.66 | 87.8 | 84.1 | 3.72 | 6.67 | 0.00 | 6.67 | 40.0 | 35.0 | 5.00 | 38.8 | 29.0 | 9.78 | 6.56 |
Qwen2.5-14B | 76.4 | 72.8 | 3.66 | 93.1 | 89.6 | 3.49 | 6.67 | 3.33 | 3.33 | 47.5 | 45.0 | 2.50 | 41.0 | 39.0 | 2.07 | 3.01 |
Qwen2.5-32B | 80.5 | 76.8 | 3.72 | 92.2 | 92.7 | -0.531 | 20.0 | 3.33 | 16.7 | 57.5 | 50.0 | 7.50 | 47.4 | 42.4 | 5.04 | 6.48 |
Gemma2-9B vs Gemma2-27B | Llama3.1-8B vs Llama3.1-70B | |||||||||||
Model | MATH | AMC | Olympiad | AIME | GSM8k | Average | MATH | AMC | Olympiad | AIME | GSM8k | Average |
Llama3.2-1B | -1.42 | -7.50 | 0.00 | 0.00 | -0.227 | -1.83 | -1.42 | -5.00 | -0.296 | 3.33 | 0.152 | -0.646 |
Llama3.2-3B | 2.08 | -7.50 | -0.888 | 0.00 | 1.67 | -0.928 | -0.14 | 10.0 | -0.593 | 3.33 | 1.06 | 2.73 |
Llama3.1-8B | 0.56 | 0.00 | 0.078 | 0.00 | -0.516 | 0.0243 | -2.18 | 7.50 | 2.67 | 0.00 | -1.29 | 1.34 |
Llama3.1-70B | 0.02 | 7.50 | -0.741 | 10.0 | 0.152 | 3.39 | 2.72 | 17.5 | 5.48 | 6.67 | 0.986 | 6.67 |
Qwen2.5-0.5B | -4.56 | 0.00 | 0.741 | 0.00 | 0.592 | -0.645 | -1.88 | 0.00 | 0.185 | 0.00 | -1.74 | -0.688 |
Qwen2.5-1.5B | -1.20 | 2.50 | -1.19 | 0.00 | -0.986 | -0.174 | -1.48 | 5.00 | -0.148 | 3.33 | -1.14 | 1.11 |
Qwen2.5-3B | 0.44 | 5.00 | 1.78 | 0.00 | -0.758 | 1.29 | -1.26 | 5.00 | -0.741 | -3.33 | -1.29 | -0.325 |
Qwen2.5-7B | 0.22 | 5.00 | 1.04 | -3.33 | 3.94 | 1.37 | 3.68 | 20.0 | 4.15 | 3.33 | 2.81 | 6.79 |
Qwen2.5-14B | 1.32 | 2.50 | -0.148 | 0.00 | -0.986 | 0.537 | 2.18 | 0.00 | 0.445 | 3.33 | -0.303 | 1.13 |
Qwen2.5-32B | 0.10 | 2.50 | 1.48 | 3.44 | 1.36 | 1.78 | 2.72 | -2.50 | 5.63 | 3.33 | 0.834 | 2.00 |
Appendix C Examples of Speaking Style Shift
We adopt the method from (Lin et al., 2023) to evaluate the most shifted tokens after fine-tuning on long CoT and Large teacher CoT data. Figure 8 shows the calculation process. This allows us to compare the token distribution shifts induced by the fine-tuning process. We annotate the tokens that exhibit the largest rank shifts as the most shifted tokens. We choose Qwen2.5-3B-Instruct as the student model. We put the results of most shifted tokens after fine-tuning on long CoT data in Figure 9 and 10. The results of most shifted tokens after fine-tuning on large teacher CoT data are shown in Figure 11. Our analysis reveals that these tokens are predominantly associated with expressive and stylistic elements, such as “wait”, “But”, and “Let”.
Appendix D Examples of Various CoT Data
This section demonstrates examples of different CoT data including long CoT, short CoT, large teacher CoT and small teacher CoT. See the following examples.